# Explainable Procedural Mistake Detection

### v1

@@ E.3 Analysis of Question Sources in In-Context Learning
Figure 6: Histograms of VLMs’ selected question sources, either explanatory dialog context or in-context learning (ICL) examples, by visual question generation (VQG) iteration for likelihood-based question selection (top) and coherence-based question selection (bottom). For a controlled comparison in this visualization, we fix the stopping criteria $n=10$, $\delta=0.1$, and $\epsilon=0.05$.
E.4 DPO Training Data Composition Ablation
In Table 7, we include an additional result for DPO where generated training data does not include candidate questions generated with in-context learning. As shown, compared to the results in Table 2, all metrics reach slightly slower maxima in this ablated configuration.
LLaVA + DPO (without ICL)
@@ E.4 DPO Training Data Composition Ablation
L 61.9 88.0 30.0 1.82 0.731
L 64.2 65.6 30.8 2.18 0.617
C 61.8 94.8 33.4 1.83 0.761
C 62.3 94.9 34.5 1.88 0.764
E.5 Question Generation Fine-Tuning and Inference Details
[…]
LLaVA + DPO
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 1e-5 0.1 8 0.1 0.05 0.50
L 6 0.05 0.025 0.38
C 6 0.1 0.05 0.04
C 4 0.05 0.05 0.08
LLaVA + DPO (without ICL)
Rank ICL $\eta$ $\beta$ $n$ $\delta$ $\epsilon$ $\tau$
L 1e-5 0.5 4 0.4 0.05 0.67
L 4 0.4 0.05 0.67
C 4 0.05 0.1 0.77
C 4 0.05 0.05 0.77
Table 8: Selected training and inference hyperparameters for the results presented in Tables 2 and 7.
LLaVA + DPO
L 64.6 94.2 30.8 1.80 0.819
L 66.4 75.5 38.4 2.65 0.759
C 61.0 95.9 27.1 1.89 0.775
C 60.8 96.7 27.1 1.79 0.768
LLaVA + DPO (without ICL)
L 62.6 87.7 29.6 1.81 0.720
L 62.8 72.0 30.6 2.07 0.651
C 65.2 94.4 34.6 1.82 0.748
C 64.8 94.5 34.4 1.88 0.747
Table 9: Ego4D-PMD validation set results for DPO-trained VLMs both with and without applying in-context learning in generating training data. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.6 GPT-4o Evaluation
To better contextualize our results with state-of-the-art proprietary large LMs, we additionally evaluated GPT-4o OpenAI et al. (2024) [^20] Specifically, we use the August 6, 2024 version of GPT-4o available in Azure AI Foundry. on our Ego4D-PMD dataset for explainable PMD. While proprietary models like GPT-4o offer limited customization, making many of the experiments we presented for open-source VLMs, evaluating vanilla GPT-4o serves as a reference point. It is worth noting that even accessed through APIs, GPT-4o returns responses too slowly to be viable for online, frame-by-frame use.
This is…from the photo. Generate an appropriate yes/no question.
Q:
We also ran into a couple issues with the Azure OpenAI API. Some specific examples in our dataset triggered Azure’s content filter, even when setting the filter’s threshold to high. To minimize this, we replaced “executed” with “complete” in the VQG prompt. In the cases where the issue persisted, we were forced to skip over the example during evaluation. Also, a small portion of API responses returned no completion content during VQG or question rephrasing for NLI (i.e. the content attribute of the response had a value of None). To deal with this, we implemented a retry procedure, where if no content is given by GPT, we prompt it again with the same prompt. If this second chance also provides no content, we either ignore the example during evaluation in the case of VQG, or simply concatenate the question and answer for question rephrasing. On the validation data, 28 out of 500 examples were omitted, and 103 question candidates could not be rephrased by GPT-4o.
The selected inference hyperparameters $n$, $\delta$, $\epsilon$, and $\tau$ are tuned as in our open-source model results, and listed in Table 10. The results on the validation data are listed in Table 11. Comparing GPT-4o to base VLMs we evaluated, it is generally outperformed open-source VLMs in PMD accuracy and informativeness, but asks more relevant questions, runs for fewer iterations, and has higher information gain. Our best models, though, outperform GPT-4o under all evaluation metrics.
$n$ $\delta$ $\epsilon$ $\tau$
4 0.4 0.1 0.19
Table 10: Selected inference hyperparameters for GPT-4o.
Validation 58.7 54.3 17.5 1.83 0.730
Table 11: Ego4D-PMD validation set results for GPT-4o.
E.7 Using Coherence Metrics to Diagnose Common VLM Behaviors
To deepen the insights from the graphs shown in Section 4.4, in Figure 7, we provide several example outputs from LLaVA with coherence-based ranking, which displays a range of behaviors. Below, we further explain these behaviors and examples.
@@ E.7 Using Coherence Metrics to Diagnose Common VLM Behaviors
Cyan points have low error with high informativeness and relevance, indicating correct decisions with coherent explanations. These are the best case examples from the model. Figure 7, Example A is one such case, where LLaVA correctly determines that the procedure “Pick up a sink brush from the kitchen slab” has been successfully completed, explaining it coherently and succinctly with a single question and answer about the location of the sink brush.
@@
Conversely, red to magenta points have high error, low informativeness, and low relevance, indicating incorrect decisions with incoherent explanations. These are the worst case examples from the model. Figure 7, Example B is one such case, where LLaVA incorrectly decides that the procedure “Grind the steel pole with the angle grinder” was not successfully completed due to the person in the image not wearing protective gear, an incoherent explanation for the decision.
@@
Indigo to black points have low error, but low relevance and informativeness, indicating correct decisions without sufficient explanation. Figure 7, Example C, is an instance of this, where LLaVA correctly decides that the person in the image has not successfully completed the procedure “Pick onion ring from the bowl of flour mixture with the fork,” explains it only by generating an answering a question about the location of a bowl of flour (an insufficient explanation).
@@
White points have high error, relevance, and informativeness, indicating coherent explanations that do not lead to a correct decision. In other words, the information collected by the VLM should have been sufficient to make a correct decision (according to our automated coherence metrics), but this did not occur. Figure 7, Example D shows one such case, where LLaVA incorrectly decides that the procedure “Drop the bottle of mustard on the countertop” was unsuccessful. While it correctly identified that the bottle is on the countertop, which suggests the success of the procedure, it later mistakenly identified the bottle to be on the floor, causing it to make the wrong decision. The ability of this analysis to easily identify issues like this may be useful for future work in PMD and task guidance, as it enables the detection and thus the correction of bugs in the system’s reasoning.
@@
Blue points have low relevance but relatively high informativeness, indicating irrelevant questions that still yield informative answers. As shown in Figure 7, Example E, this does not necessarily indicate a failure of LLaVA, rather a terse explanation. In this example, LLaVA correctly determines that the procedure “Pick the manual paper” has not been completed successfully. It reasonably explains this decision by asking about the presence of a manual paper and responding with No. The question of whether the person is working with a manual paper is deemed irrelevant by our metrics because if the answer were instead Yes, this would not provide sufficient information to conclude that the procedure was successful. However, since the answer was No, we do have sufficient information to conclude that the procedure is unsuccessful, despite the question being relatively indirect. Blue points may thus point to sufficient explanations which lack some detail or specificity.
@@
Green and yellow points have high relevance but low informativeness, indicating a failure to extract useful information in VQA. Green points have close to zero informativeness, indicating unsure responses in VQA. In Figure 7, Example F, LLaVA explains its decision about the procedure “Stir the chocolate inside the blender” by asking whether the chocolate and the spoon are in the blender in various ways. However, these objects are not present in the image and thus LLaVA’s answer is not confident, causing it to respond Unsure to most questions, leading to low informativeness. Despite its failure to answer questions, LLaVA still arrives at the correct conclusion that the procedure has not been successfully completed.
@@
Example A: Pick up a sink brush from the kitchen slab. Refer to caption Label: ✓        Predicted: ✓ Explanation: 1. Is the sink brush in the person’s hands? Yes
Example B: Grind the steel pole with the angle grinder. Refer to caption Label: ✓        Predicted: ✗ Explanation: 1. Is the person wearing safety goggles? No 2. Is the person wearing any protective gear? No 3. Is the person wearing gloves? No
Example C: Pick onion ring from the bowl of flour mixture with the fork. Refer to caption Label: ✗        Predicted: ✗ Explanation: 1. Is the bowl of flour mixture on the table? Yes
Example D: Drop the bottle of mustard on the countertop. Refer to caption Label: ✓        Predicted: ✗ Explanation: 1. Is the bottle of mustard on the countertop? Yes 2. Is the bottle of mustard on the floor? Yes
Example E: Pick the manual paper. Refer to caption Label: ✗        Predicted: ✗ Explanation: 1. Is the person working on a manual paper? No
Example F: Stir the chocolate inside the blender with the spoon. Refer to caption Label: ✗        Predicted: ✗ Explanation: 1. Is the chocolate in the blender? Unsure 2. Is the chocolate inside the blender? Yes 3. Is the spoon in the blender? Unsure 4. Is the spoon in the chocolate? Unsure
Example G: Put the trowel in a bin. Refer to caption Label: ✓        Predicted: ✗ Explanation: 1. Is the trowel in a bin? No
Example H: Put the bottle in the cabinet. Refer to caption Label: ✗        Predicted: ✓ Explanation: 1. Is the bottle in the cabinet? Yes
Figure 7: Sample outputs from LLaVA with coherence-based ranking, representing the range of behaviors visualized in Figure 4. Some images cropped for clarity and space.

### v2

@@ E.7 Analysis of Question Sources in In-Context Learning
Figure 6: Histograms of VLMs’ selected question sources, either self-dialog context or in-context learning (ICL) examples, by visual question generation (VQG) iteration for likelihood-based question selection (top) and coherence-based question selection (bottom).
E.8 Diversity-Based Ranking Baseline
One possible explanation for the effectiveness of coherence-based question ranking in Section 4.1 is that it enables the selection of more semantically diverse questions, thus collecting broader information about the image. To explore this question, we implement a supplemental diversity-based ranking approach which uses a sentence transformer Reimers and Gurevych (2020) [^25] See https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2. to embed all previous and candidate questions at each iteration, then selects the candidate question with the largest average cosine distance from previous questions.
As shown in Table 16, we observe that diversity-based ranking combined with in-context learning can also improve the accuracy of VLMs (to a level slightly below that of coherence-based ranking). Accuracy reaches respective maxima of 64.7% and 67.1% for InstructBLIP and LLaVA (compared to 66.6% and 67.8% under coherence-based ranking). This may suggest that some accuracy improvements in coherence-based ranking could be attributed simply to the ability to select more diverse questions than a likelihood-based approach. Additioanlly, this reaffirms our observation that VLMs are poorly suited for this task off-the-shelf.
However, we also see that in most cases, Diversity-based ranking substantially degrades relevance, informativeness, number of iterations, and information gain compared to those in Table 2. This suggests that the highly exploratory nature of diversity-based ranking causes rationales to be less coherent, and conclusions are made slower and less confidently with this approach. Meanwhile, coherence-based ranking enables us to find good questions to ask faster, leading to more confident conclusions with more relevant and informative supporting evidence (while also achieving a higher accuracy).
InstructBLIP
ICL $n$ $\delta$ $\epsilon$ $\tau$
10 0.1 0.2 0.25
10 0.1 0.1 0.23
LLaVA
ICL $n$ $\delta$ $\epsilon$ $\tau$
10 0.05 0.05 0.58
10 0.05 0.05 0.69
Llama 3
ICL $n$ $\delta$ $\epsilon$ $\tau$
10 0.05 0.05 0.41
10 0.1 0.025 0.40
Table 15: Selected inference hyperparameters for diversity-based ranking with various VLMs, with optional supplementary candidates generated through in-context learning (ICL).
InstructBLIP
Partition ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 61.4 17.8 .244 2.89 .268
Validation 64.0 14.4 .337 3.82 .350
Testing 64.2 17.5 .237 2.88 .268
Testing 64.7 15.1 .337 3.72 .344
LLaVA
Partition ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 64.2 37.0 .298 4.12 .444
Validation 65.2 43.2 .385 3.99 .539
Testing 62.9 35.2 .310 4.35 .455
Testing 67.1 39.8 .405 4.32 .513
Llama 3
Partition ICL Acc. $\uparrow$ Rel. $\uparrow$ Inf. $\uparrow$ # Iter. $\downarrow$ I. Gain $\uparrow$
Validation 62.6 15.7 .313 6.19 .262
Validation 66.8 24.2 .412 5.14 .358
Testing 60.9 15.1 .322 6.48 .246
Testing 59.6 22.6 .393 5.23 .332
Table 16: Ego4D-PMD validation and testing set results for diversity-based ranking with various VLMs, with optional supplementary candidates generated through in-context learning (ICL).
E.9 DPO with Length Penalty Experiment
In the results discussed in Section 5.2, we observed that while VLMs learned to generate much more relevant questions, the informativeness of answers and thus the PMD accuracy degraded. We hypothesized that this resulted from the generation of highly complex questions, e.g., “Is the soil placed around the seedling with the trowel in the person’s hand?” As such, a potentially fruitful avenue for future research is to explore decoding approaches and learning objectives that prioritize more approachable questions for VLMs.
While the primary purpose of this work was to recast the problem of PMD into a more transparent formulation and lay a foundation for research toward coherent question generation and answering in PMD, we performed an initial experiment to guide future work along this line. Specifically, we imposed an exponential length penalty $l=-1.0$ to text generation during training data generation and inference for evaluation. [^26] Like the results presented in Table 2, candidates generated through in-context learning are also included in training data. During beam search, a length penalty modifies the total log-likelihood $p_{Q}$ of a partial candidate question $Q$ as follows:
$p^{\prime}_{Q}=\frac{p_{Q}}{{|Q|}^{l}}$
Here, $|Q|$ is the length of the question $Q$, i.e., number of generated tokens thus far. For earlier experiments, the value of $l$ could be thought of as 1, which effectively applies no exponential penalty to $|Q|$ when calculating likelihood-based sequence scores (the default behavior).
In Table 17, we list the results of applying this length penalty. As shown, compared to the results in Table 2, similar trends of hold despite applying the length penalty: we observe improved relevance, number of iterations, and information gain, but degraded accuracy and informativeness. An interesting exception is when using likelihood-based ranking and in-context learning during inference, we recover a comparable accuracy and informativeness to those observed in LLaVA before applying DPO, but this comes at a cost of a a lower relevance, higher number of iterations, and lower information gain than other inference configurations. This provides further evidence that there exists a trade-off between generating relevant questions and achieving high informativeness and accuracy, and future work should aim to find a balance between these priorities.
LLaVA + DPO (with length penalty)
@@ E.9 DPO with Length Penalty Experiment
L 61.7 93.5 .293 1.93 .741
L 67.0 56.4 .459 4.05 .628
C 63.1 97.5 .313 1.82 .771
C 62.9 97.7 .319 1.83 .769
Table 17: Ego4D-PMD test set results for DPO-trained VLMs with an additional length penalty $l=-1$ applied during training data generation and inference. Inference applies likelihood (L) or coherence (C) candidate question ranking approaches, with optional supplementary candidates generated through in-context learning (ICL).
E.10 Using Coherence Metrics to Diagnose Common VLM Behaviors
To deepen the insights from the graphs in Figure 4, in Figure 7, we provide several example outputs from LLaVA with coherence-based ranking, which displays a range of behaviors. Below, we further explain these behaviors and examples.
@@ E.10 Using Coherence Metrics to Diagnose Common VLM Behaviors
Cyan points have low error with high informativeness and relevance, indicating correct decisions with coherent rationales. These are the best case examples from the model. Figure 7, Example A is one such case, where LLaVA correctly determines that the procedure “Pick up a sink brush from the kitchen slab” has been successfully completed, rationalizing it coherently and succinctly with a single question and answer about the location of the sink brush.
@@
Conversely, red to magenta points have high error, low informativeness, and low relevance, indicating incorrect decisions with incoherent rationales. These are the worst case examples from the model. Figure 7, Example B is one such case, where LLaVA incorrectly decides that the procedure “Tighten the screw” was not successfully completed due to the person in the image not wearing various protective gear, an incoherent rationale for the decision.
@@
Indigo to black points have low error, but low relevance and informativeness, indicating correct decisions without sufficient rationale. Figure 7, Example C, is an instance of this, where LLaVA correctly decides that the person in the image has not successfully completed the procedure “Paint the stone” (rather they are painting a wood molding). However, LLaVA’s decision is only supported by a question about whether the person is wearing a shirt, which it does not answer confidently, making for a completely insufficient rationale.
@@
White points have high error, relevance, and informativeness, indicating coherent rationales that do not lead to a correct decision. In other words, the information collected by the VLM should have been sufficient to make a correct decision (according to our automated coherence metrics), but this did not occur. Figure 7, Example D shows one such case, where LLaVA incorrectly decides that the procedure “Drop the bottle of mustard on the countertop” was unsuccessful. While it correctly identified that the bottle is on the countertop, which suggests the success of the procedure, it later mistakenly identified the bottle to be on the floor, creating a contradiction in the rationale and causing it to make the wrong decision. The ability of this analysis to easily identify issues like this may be useful for future work in PMD and task guidance, as it enables the detection and thus the correction of bugs in the system’s reasoning.
@@
Blue points have low relevance but relatively high informativeness, indicating irrelevant questions that still yield informative answers. As shown in Figure 7, Example E, this does not necessarily indicate a failure of LLaVA, rather a terse rationale. In this example, LLaVA correctly determines that the procedure “Fold the cut piece of cloth” has not been completed successfully. It reasonably rationalizes this decision by asking about the presence of a piece of cloth and responding with No. The question of whether the person is working with a piece of cloth is deemed somewhat irrelevant by our metrics because if the answer were instead Yes, this would not provide sufficient information to conclude that the procedure was successful. However, since the answer was No, we do have sufficient information to conclude that the procedure is unsuccessful, despite the question being relatively indirect. Blue points may thus point to sufficient rationales which lack some detail or specificity, which are not necessarily problematic to system performance.
@@
Green and yellow points have high relevance but low informativeness, indicating a failure to extract useful information in VQA. Green points have close to zero informativeness, indicating unsure responses in VQA. In Figure 7, Example F, LLaVA rationalizes its decision about the procedure “Put the bolt remover in the lawn tractor” by asking whether the bolt remover is in the lawn tractor in various ways. However, these objects are not present in the image and thus LLaVA’s answer is not confident, causing it to respond Unsure to most questions, leading to low informativeness. Despite its failure to answer questions, LLaVA still arrives at the correct conclusion that the procedure has not been successfully completed.
@@
The colors of these points can be used to characterize common behaviors of VLMs. Additional insights toward the fine-grained strengths and weaknesses of various approaches may be gained from analyzing these results by mistake type, or verb and noun categories.
Example A: Pick up a sink brush from the kitchen slab. Refer to caption Label: ✓        Predicted: ✓ Rationale: 1. Is the sink brush in the person’s hand? Yes
Example B: Tighten the screw. Refer to caption Label: ✓        Predicted: ✗ Rationale: 1. Is the person wearing gloves? No 2. Is the person wearing protective gear? No 3. Is the person wearing a mask? No
Example C: Paint the stone. Refer to caption Label: ✗        Predicted: ✗ Rationale: 1. Is the person wearing a shirt? Unsure
Example D: Drop the bottle of mustard on the countertop. Refer to caption Label: ✓        Predicted: ✗ Rationale: 1. Is the bottle of mustard on the countertop? Yes 2. Is the bottle of mustard on the floor? Yes
Example E: Fold the cut piece of cloth. Refer to caption Label: ✗        Predicted: ✗ Rationale: 1. Is the person working on a piece of cloth? No
Example F: Put the bolt remover in the lawn tractor. Refer to caption Label: ✗        Predicted: ✗ Rationale: 1. Is the bolt remover in the lawn tractor? Unsure 2. Is the bolt remover inside the lawn tractor? Unsure 3. Is the bolt remover attached to the lawn tractor? Unsure
Example G: Put the trowel in a bin. Refer to caption Label: ✓        Predicted: ✗ Rationale: 1. Is the trowel in a bin? No
Example H: Put the bottle in the cabinet. Refer to caption Label: ✗        Predicted: ✓ Rationale: 1. Is the bottle in the cabinet? Yes
Figure 7: Sample outputs from LLaVA with coherence-based ranking, representing the range of behaviors visualized in Figure 4. Images cropped for visibility and space.